\begin{thebibliography}{}	
				
		\bibitem[Alonso 2014]{Alonso2014}Alonso LM, Proekt A, Schwartz TH, Pryor KO, Cecchi GA, Magnasco MO. Dynamical criticality during induction of anesthesia in human ECoG recordings. Frontiers in neural circuits. 2014 Mar 25;8:20.
		
		\bibitem[Anderson 2010]{Anderson2010}Anderson GW, Guionnet A, Zeitouni O. An introduction to random matrices, volume 118 of Cambridge Studies in Advanced Mathematics.
		
		\bibitem[Andronov 2013]{Andronov2013} Andronov AA, Vitt AA, Khaikin SE. Theory of Oscillators: Adiwes International Series in Physics. Elsevier; 2013 Oct 22.		
		
		\bibitem[Bienenstock 1998]{bienenstock1998regulated} Bienenstock E, Lehmann D. Regulated criticality in the brain?. Advances in complex systems. 1998 Dec;1(04):361-84.	

		\bibitem[Camalet 2000] {camalet2000auditory} Camalet S, Duke T, Julicher F, Prost J. Auditory sensitivity provided by self-tuned critical oscillations of hair cells. Proceedings of the National Academy of Sciences. 2000 Mar 28;97(7):3183-8. 
		
		\bibitem[Choe 1998]{choe1998model} Choe Y, Magnasco MO, Hudspeth AJ. A model for amplification of hair-bundle motion by cyclical binding of Ca2+ to mechanoelectrical-transduction channels. Proceedings of the National Academy of Sciences. 1998 Dec 22;95(26):15321-6.
		
		\bibitem[Churchland 2012]{churchland2012} Churchland MM, Cunningham JP, Kaufman MT, Foster JD, Nuyujukian P, Ryu SI, Shenoy KV. Neural population dynamics during reaching. Nature. 2012 Jul;487(7405):51.
		
		\bibitem[Duke 2003]{Duke2003} Duke T, Julicher F. Active traveling wave in the cochlea. Physical review letters. 2003 Apr 16;90(15):158101.
		
		\bibitem[Egu\'{i}luz 2000] {eguiluz2000essential} Egu\'{i}luz VM, Ospeck M, Choe Y, Hudspeth AJ, Magnasco MO. Essential nonlinearities in hearing. Physical Review Letters. 2000 May 29;84(22):5232.
		
		\bibitem[Freeman 2005]{freeman2005metastability} Freeman WJ, Holmes MD. Metastability, instability, and state transition in neocortex. Neural Networks. 2005 Aug 31;18(5):497-504.
		
		\bibitem[Freire 1988]{Freire1988Algorithm} Freire E, Gamero E, Ponce E, Franquelo LG. An algorithm for symbolic computation of center manifolds. InInternational Symposium on Symbolic and Algebraic Computation 1988 Jul 4 (pp. 218-230). Springer, Berlin, Heidelberg.
		
		\bibitem[Gilbert 2007] {gilbert2007} Gilbert CD, Sigman M. Brain states: top-down influences in sensory processing. Neuron. 2007 Jun 7;54(5):677-96.
		
		\bibitem[Gilbert 2013] {gilbert2013} Gilbert CD, Li W. Top-down influences on visual processing. Nature Reviews Neuroscience. 2013 May;14(5):350.
		
		\bibitem[Hayton 2018a]{Hayton2018a} Hayton K, Moirogiannis D, Magnasco M (2018) Adaptive scales of integration and response latencies in a critically-balanced model of the primary visual cortex. PLoS ONE 13(4): e0196566.
		
		\bibitem[Hayton 2018b]{Hayton2018b} Hayton K, Moirogiannis D, Magnasco M. A Nonhyperbolic Toy Model of Cochlear Dynamics. arXiv preprint arXiv:1807.00764. 2018 Jul 2.
		
		\bibitem[Hopf 1942]{Hopf1942}Hopf E. Abzweigung einer periodischen Lösung von einer stationären Lösung eines Differentialsystems. Ber. Math.-Phys. Kl Sächs. Akad. Wiss. Leipzig. 1942 Jan;94:1-22.
		
		\bibitem[Hoppensteadt 2012]{hoppensteadt2012weakly} Hoppensteadt FC, Izhikevich EM. Weakly connected neural networks. Springer Science and Business Media; 2012 Dec 6.
		
		\bibitem[Ioos 1999]{Ioos1999} Ioos G,  Adelmeyer M.  Topics in Bifurcation Theory and Applications. Advanced Series in Nonlinear Dynamics. World Scientific, 2nd edition, 1999
		
		\bibitem[Kapadia 1999] {kapadia1999dynamics} Kapadia MK, Westheimer G, Gilbert CD. Dynamics of spatial summation in primary visual cortex of alert monkeys. Proceedings of the National Academy of Sciences. 1999 Oct 12;96(21):12073-8.
		
		\bibitem[Kern 2003] {kern2003essential} Kern A, Stoop R. Essential role of couplings between hearing nonlinearities. Physical Review Letters. 2003 Sep 19;91(12):128101.
		
		\bibitem[Kirst 2017] {kirst2017} Kirst C, Modes CD, Magnasco MO. Shifting attention to dynamics: Self-reconfiguration of neural networks. Current Opinion in Systems Biology. 2017 Jun 1;3:132-40.

		\bibitem[Kuznetsov 2013]{Kuznetsov2013} Kuznetsov YA. Elements of applied bifurcation theory. Springer Science and Business Media; 2013 Mar 9.

		\bibitem[Machens 2005]{machens2005} Machens CK, Romo R, Brody CD. Flexible control of mutual inhibition: a neural model of two-interval discrimination. Science. 2005 Feb 18;307(5712):1121-4.
		
		\bibitem[Magnasco 2003]{Magnasco2003} Magnasco MO. A wave traveling over a Hopf instability shapes the cochlear tuning curve. Physical review letters. 2003 Feb 4;90(5):058101.

		\bibitem[Moirogiannis 2017]{Moirogiannis2017} Moirogiannis D, Piro O, Magnasco MO. Renormalization of Collective Modes in Large-Scale Neural Dynamics. Journal of Statistical Physics. 2017 May 1;167(3-4):543-58.
		
		\bibitem[Moirogiannis 2019]{Moirogiannis2019} Moirogiannis D, Hayton K, Magnasco M. A Center Manifold Reduction Technique for a System of Randomly Coupled Oscillators. arXiv preprint arXiv:1904.00892. 2019 Apr 1.
		
		\bibitem[Nauhaus 2009]{nauhaus2009stimulus} Nauhaus I, Busse L, Carandini M, Ringach DL. Stimulus contrast modulates functional connectivity in visual cortex. Nature neuroscience. 2009 Jan 1;12(1):70-6.

		\bibitem[O'Rourke 2016]{O'Rourke2016Eigenvectors}O'Rourke S, Vu V, Wang K. Eigenvectors of random matrices: a survey. Journal of Combinatorial Theory, Series A. 2016 Nov 1;144:361-442.
		
		\bibitem[Poincar\`{e} 1893]{poincare1893}Poincar\`{e} H. Les méthodes nouvelles de la mécanique céleste: Méthodes de MM. Newcomb, Glydén, Lindstedt et Bohlin. 1893. Gauthier-Villars it fils; 1893.
		
		\bibitem[Sceniak 1999] {sceniak1999contrast} Sceniak MP, Ringach DL, Hawken MJ, Shapley R. Contrast's effect on spatial summation by macaque V1 neurons. Nature neuroscience. 1999 Aug 1;2(8):733-9.

		\bibitem[Seung 1998]{seung1998continuous} Seung HS. Continuous attractors and oculomotor control. Neural Networks. 1998 Nov 30;11(7):1253-8.
		
		\bibitem[Seung 2000]{seung2000stability} Seung HS, Lee DD, Reis BY, Tank DW. Stability of the memory of eye position in a recurrent network of conductance-based model neurons. Neuron. 2000 Apr 30;26(1):259-71.

		\bibitem[Solovey 2012]{Solovey2012}Solovey G, Miller KJ, Ojemann J, Magnasco MO, Cecchi GA. Self-regulated dynamical criticality in human ECoG. Frontiers in integrative neuroscience. 2012 Jul 19;6:44.
		
		\bibitem[Solovey 2015]{solovey2015loss} Solovey G, Alonso LM, Yanagawa T, Fujii N, Magnasco MO, Cecchi GA, Proekt A. Loss of consciousness is associated with stabilization of cortical activity. Journal of Neuroscience. 2015 Jul 29;35(30):10866-77.

		\bibitem[Tao 2008]{Tao2008circular}Tao T, Vu V. Random matrices: the circular law. Communications in Contemporary Mathematics. 2008 Apr;10(02):261-307.
		
		\bibitem[Tao 2012]{Tao2012Random}Tao T, Vu V. Random matrices: Universal properties of eigenvectors. Random Matrices: Theory and Applications. 2012 Jan;1(01):1150001.

		\bibitem[Wiggins 2003]{wiggins2003introduction} Wiggins S. Introduction to applied nonlinear dynamical systems and chaos. Springer Science and Business Media; 2003 Oct 1.
		
		\bibitem[Yan 2012]{yan2012input} Yan XH, Magnasco MO. Input-dependent wave attenuation in a critically-balanced model of cortex. PloS one. 2012 Jul 25;7(7):e41419.
		
		
	\end{thebibliography}

\documentclass[english, aps, showpacs, preprint, nobibnotes]{revtex4-1}

\usepackage{amsmath,amssymb}

\usepackage[utf8x]{inputenc}

\usepackage[T1]{fontenc}
\usepackage{float}
\usepackage{mathtools}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{graphicx}
\usepackage{nicefrac}
\usepackage{resizegather}
\usepackage{breakcites}

\begin{document}
	
	\title{Emergence of Subcritical Bifurcations in a System of Randomly Coupled Supercritical Andronov-Hopf Oscillators: A Potential Mechanism for Neural Network Type Switching}
	\date{\today}
	
	\author{Keith Hayton}
	\email{khayt86@gmail.com }
	\thanks{author contributed equally}
	\author{Dimitrios Moirogiannis}
	\email{dmoirogi@gmail.com}
	\thanks{author contributed equally}
	\affiliation{Center for Studies in Physics and Biology, The Rockefeller University}

	\begin{abstract}
		
	Experimental evidence suggests that the computational state of cortical systems change according to behavioral and stimulus context. However, it is still unknown what mechanisms underlie this adaptive processing in cortical circuitry. In this paper, we present a model of randomly coupled supercritical Andronov-Hopf oscillators which can act as an adaptive processor by exhibiting drastically different dynamics depending on the value of a single network parameter. Despite being only composed of supercritical subunits, the full system can exhibit either supercritical or subcritical Andronov-Hopf bifurcations. This model might provide a novel mechanism for switching between globally asymptotically stable and nonhyperbolic neural network types in pattern recognition theory.
		
	\end{abstract}

	\maketitle
	
	\section*{Introduction}
	
	A growing body of evidence suggests that cortical areas are not fixed in function but instead act as adaptive processors \cite{kirst2017} with intrinsic cortical circuits that are strongly modulated by feedback from higher cortical areas \cite{gilbert2007, gilbert2013} and incoming sensory information \cite{kapadia1999dynamics, sceniak1999contrast, nauhaus2009stimulus, yan2012input, Hayton2018a}. This is in contrast to the classical feedforward view of the cortex as a hierarchical series of cortical areas which encode an increasingly more complex description of sensory input \cite{gilbert2007, gilbert2013}. How the brain is able to dynamically reconfigure it's computational state is largely unknown and a matter of active research \cite{kirst2017}.
	
	In this paper, we present a toy model of cortical circuitry consisting of a network of damped, identical oscillators coupled through a \textit {Wigner real symmetric matrix} \cite{Tao2008circular, Anderson2010, Tao2012Random, O'Rourke2016Eigenvectors}. Each oscillator in the network is poised at a supercritical \textit{Andronov-Hopf bifurcation} \cite{poincare1893, Hopf1942, Andronov2013}. To study the dynamics \cite{Ioos1999, wiggins2003introduction, Kuznetsov2013} of the system, we follow the statistical \textit{center manifold reduction} approach introduced in \cite{Moirogiannis2019}. We show that by changing a single network parameter the full system can undergo either a supercritical or subcritical Andronov-Hopf bifurcation. The transformation between supercritical and subcritical occurs despite the system being exclusively composed of supercritical oscillators. This model is an example of how a network of fixed computational subunits can act as an adaptive processor by drastically switching it's dynamic regime without affecting the subunit dynamics. Additionally, we also explore how this network model might also act as a neuroscience inspired adaptive pattern recognition processor by switching between neural network types, specifically \textit{globally asymptotically stable} and \textit{nonhyperbolic}. 
	
	\section*{Analysis}
	
	We consider the following system of $N$ coupled, damped, identical oscillators:
	
	\begin{eqnarray}
	\dot{x}_{i}&=&y_{i}+ax_{i}^2+bx_{i}^3+\underset{j}{\sum}M_{ij}x_{j}\label{eq:system}\\
	\dot{y}_{i}&=&-x_{i}\nonumber 
	\end{eqnarray}
	
	\noindent , where $i\in\left\{ 1,\ldots, N\right\} $ and $a,b \in \mathbb{R}$ are system parameters. The individual subunits
	
	\begin{eqnarray}
	\dot{x}_{i}&=&{y}_{i}+a{x}_{i}^2+b{x}_{i}^3\label{eq:units}\\
	\dot{y}_{i}&=&-x_{i}\nonumber 
	\end{eqnarray}	
	
	\noindent are linearly coupled through the variables $x_{i}$, with the connectivity matrix $M \in \mathbb{R}^{N \times N}$ given by a random matrix with all eigenvalues negative and contained within a compact set in the complex plane except for a single leading eigenvalue $\lambda$, which will act as a bifurcation parameter. The construction of such a matrix is outlined in \cite{Moirogiannis2019}. 
    
    In general, a system 
    
    \begin{eqnarray}
    	\dot{x}&=& y+f\left(x,y\right)\label{eq:Hopf}\\
    	\dot{y}&=&-x+g\left(x,y\right)\nonumber
    \end{eqnarray}
    	
    \noindent is poised at an Andronov-Hopf bifurcation if $f$ and $g$, which capture all nonlinearities of order $\geq2$, ensure that the first Lyaponuv coefficient $l_1(0)=\dfrac{1}{16}\left(f_{xxx}+f_{xyy}+g_{xxy}+g_{yyy}\right)+\dfrac{1}{16}\left(f_{xy}\left(f_{xx}+f_{yy}\right)-g_{xy}\left(g_{xx}+g_{yy}\right)-f_{xx}g_{xx}+f_{yy}g_{yy}\right) \neq 0$ \cite{{Kuznetsov2013}}. The bifurcation is subcritical (supercritical) iff $l_1(0) > 0$ ($l_1(0) < 0$).
    
    For the individual subunits described in (\ref{eq:units}), the first Lyapunov coefficient is given by $l_1(0)=\dfrac{1}{16}\left(f_{xxx}+f_{xyy}\right)+\dfrac{1}{16}\left(f_{xy}\left(f_{xx}+f_{yy}\right)\right)=\dfrac{1}{2}b$, where $b$ is the coefficient of the cubic term in (\ref{eq:units}). Therefore, the type of bifurcation that a single subunit is poised at is determined exclusively by $b$, with the bifurcation being subcritical (supercritical) iff $b>0$ ($b<0$); the coefficient $a$ of $x^2$ in (\ref{eq:units}) does not play a role.
    
	We will now proceed to calculate the Lyapunov coefficient for the full system. Using the statistical center manifold reduction technique outlined in \cite{Moirogiannis2019}, it has been shown that at the bifurcation point $\lambda=0$, the system of equations in (\ref{eq:system}) can be reduced to a set of restricted equations on the center manifold. Up to third order, these reduced equations are given by:
	
	\begin{eqnarray}
		\nonumber\dot{x}&=&y+a\sqrt{N}\Gamma_{1}^{2\delta_{1,\bullet}}x^{2}+\left(b{\sqrt{N}}^{2}\Gamma_{1}^{3\delta_{1,\bullet}}+2a{\sqrt{N}}^{2}\underset{k=2}{\overset{N}{\sum}}\Gamma_{1}^{\delta_{1,\bullet}+\delta_{k,\bullet}}\alpha_{0}^{k}\right)x^{3}+\\ 
		 +&&\left(2a{\sqrt{N}}^{2}\underset{k=2}{\overset{N}{\sum}}\Gamma_{1}^{\delta_{1,\bullet}+\delta_{k,\bullet}}\alpha_{1}^{k}\right)x^{2}y+\left(2a{\sqrt{N}}^{2}\underset{k=2}{\overset{N}{\sum}}\Gamma_{1}^{\delta_{1,\bullet}+\delta_{k,\bullet}}\alpha_{2}^{k}\right)xy^{2}+O\left(4\right)\\ 
		\nonumber \dot{y}&=&-x
	\end{eqnarray}
	
	\noindent , where the $\Gamma$'s are random variables which depend on the structure of the eigenvectors and are defined in \cite{Moirogiannis2019}. In the limit $N \rightarrow \infty$, the $\Gamma$'s can be evaluated using moments of the statistical distribution of the eigenvectors \cite{O'Rourke2016Eigenvectors, Moirogiannis2019}.

	By analyzing the reduced equations, it can be proved that the system undergoes an Andronov-Hopf bifurcation when $\lambda=0$. Observe that the reduced equations are of the form (\ref{eq:Hopf}), and thus, the Lyapunov coefficient can be easily calculated: $l_1(0)=b\dfrac{3}{8}{\sqrt{N}}^{2}\Gamma_{1}^{3\delta_{1,\bullet}}+a^{2}{\sqrt{N}}^{2}\underset{k=2}{\overset{N}{\sum}}\Gamma_{1}^{\delta_{1,\bullet}+\delta_{k,\bullet}}\Gamma_{k}^{2\delta_{1,\bullet}}\dfrac{-\lambda_{k}}{4\lambda_{k}^{2}+9}\neq0$. As $l_1(0)$ only vanishes along a single parabolic curve in the $a,b$ parameter plane, this implies that the system typically undergoes a Andronov-Hopf bifurcation as $\lambda$ crosses the imaginary axis. The bifurcation is subcritical (supercritical) iff $l_1(0)>0$ ($l_1(0)<0$). 
	
	Moreover, assuming the connection matrix $M$ is normal, it can be easily shown that the term ${\sqrt{N}}^{2}\underset{k=2}{\overset{N}{\sum}}\Gamma_{1}^{\delta_{1,\bullet}+\delta_{k,\bullet}}\Gamma_{k}^{2\delta_{1,\bullet}}\dfrac{-\lambda_{k}}{4\lambda_{k}^{2}+9}$ is typically positive. First, since $M$ is normal, the $\Gamma$'s can be greatly simplified: $\Gamma_{1}^{\delta_{1,\bullet}+\delta_{k,\bullet}}=\Gamma_{k}^{2\delta_{1,\bullet}}=\underset{\phi}{\sum}V_{\phi,k}V_{\phi,1}^2$ where $(V_{1,k},\dots ,V_{N,k})^t$ is the $k$th vector in the orthonormal base of eigenvectors (with eigenvalue $\lambda_{k}$) \cite{Moirogiannis2019}. It then follows that ${\sqrt{N}}^{2}\underset{k=2}{\overset{N}{\sum}}\Gamma_{1}^{\delta_{1,\bullet}+\delta_{k,\bullet}}\Gamma_{k}^{2\delta_{1,\bullet}}\dfrac{-\lambda_{k}}{4\lambda_{k}^{2}+9}={\sqrt{N}}^{2}\underset{k=2}{\overset{N}{\sum}}(\underset{\phi}{\sum}V_{\phi,k}V_{\phi,1}^2)^2\dfrac{-\lambda_{k}}{4\lambda_{k}^{2}+9}> 0$ almost surely. We also have that the other term in the Lyapunov coefficient is positive $\Gamma_{k}^{3\delta_{1,\bullet}}>0$, since $\Gamma_{k}^{3\delta_{1,\bullet}}=\underset{\phi}{\sum}V_{\phi,k}V_{\phi,1}^3=\underset{\phi}{\sum}V_{\phi,1}^4>0$.
	
	We now examine how the bifurcation of the full system and individual subunits vary with parameters $a$ and $b$. If $b>0$, the full system along with each subunit ($l_1(0)_{subunit}=\dfrac{1}{2}b$) is subcritical regardless of the value of $a$. When $b<0$, the subunits are always supercritical, but the full system can be either subcritical or supercritical depending on the value of $||a||$: the full system is subcritical (supercritical) iff $||a||>\gamma \sqrt{-b}$ ($||a||<\gamma \sqrt{-b}$), where $\gamma=\sqrt{\nicefrac{\frac{3}{8}\underset{\phi}{\sum}V_{\phi,1}^4} {\left(\underset{k=2}{\overset{N}{\sum}}(\underset{\phi}{\sum}V_{\phi,k}V_{\phi,1}^2)^2\frac{-\lambda_{k}}{4\lambda_{k}^{2}+9}\right)}}$ . Thus, it is possible for the full system to undergo a subcritical Andronov-Hopf bifurcation despite each individual subunit being supercritical.  
	
	\section*{Adaptive Pattern Recognition}
	
	We now explain how our results might be applicable to pattern recognition and switching of neural network types. 	Assuming static input patterns, neural networks can be broadly categorized into three types: \textit{multiple attractor} (MA), \textit{globally asymptotically stable} (GAS), and \textit{nonhyperbolic} (NH) neural networks \cite{hoppensteadt2012weakly}. MA-type networks, which include Hopfield and Cohen-Grossberg networks, are characterized by each input pattern converging to one of many multiple attractors; each attractor corresponds to a different stored memory pattern. GAS-type networks, on the other hand, store memories by associating a given input pattern with the globally asymptotically stable state of the system. In this case, a memory is stored as a specific location of the attractor, and an input pattern is presented to the network as a parameter which controls the attractor location \cite{hoppensteadt2012weakly}. 
	
	Finally, NH-types have nonhyperbolic equilibriums which affect the global dynamics of the system. Input in NH types is given to the system in the form of a bifurcation parameter that perturbs the nonhyperbolic equilibrium, causing it to lose stability or disappear altogether. Only a local analysis near the equilibrium is needed to understand global behavior, since the equilibrium is a point of convergence of the curves, called separatrices, separating the different attraction domains of the system \cite{hoppensteadt2012weakly}. Nonhyperbolic equilibrium dynamics have been the focus of an extensive number of studies in neuroscience. These studies include entire hemisphere ECoG recordings \cite{Solovey2012, Alonso2014, solovey2015loss}, experimental studies in premotor and motor cortex \cite{churchland2012}, theoretical \cite{seung1998continuous} and experimental studies \cite{seung2000stability} of \textit{slow manifolds} (a specific case of center manifolds) in oculomotor control, slow manifolds in decision making \cite{machens2005}, Andronov-Hopf bifurcation \cite{poincare1893, Hopf1942, Andronov2013} in the olfactory system \cite{freeman2005metastability} and cochlea \cite{choe1998model, eguiluz2000essential, camalet2000auditory, kern2003essential, Duke2003, Magnasco2003, Hayton2018b}, a nonhyperbolic model of primary visual cortex (V1) \cite{yan2012input, Hayton2018a}, and theoretical work on regulated criticality \cite{bienenstock1998regulated}.  
	
 	Following the lead of \cite{hoppensteadt2012weakly}, we consider a specific case of (\ref{eq:system}), where the $N$ damped oscillators in the system, given by (\ref{eq:units}), are neural relaxation oscillators with input matrix $R=\begin{pmatrix}r_{1} & & 0 \\ & \ddots\\ 0 & & r_{N} \end{pmatrix}$, where $r_{i}$ is the external input to the $i$-th oscillator, and $M=R+C$, where $C$ represents the linear connectivity between the $x_{i}$'s. In the case of neural relaxation oscillators, $x_{i}$ and $y_{i}$ represent rescaled populations of excitatory and inhibitory populations, respectively. As in \cite{hoppensteadt2012weakly}, we assume Dales principle holds, $C_{ij}{\geq0}$ ${\forall}i,j$, and because of Hebbian learning mechanisms, $C$ is symmetric.
	
	It has been shown, using the Perron-Frobenius theorem, that for any input matrix $R$, the leading eigenvalue $\lambda$ of $R+C$ is simple, and if $\lambda=0$, the full system undergoes an Andronov-Hopf bifurcation \cite{hoppensteadt2012weakly}. Using this result, it has also been shown that such a system has the potential to perform pattern recognition on a set of input vectors. The exact construction of the pattern recognition algorithm is still an open problem as it requires the construction of a specific matrix for all input vectors, however, once this matrix is constructed, the resulting algorithm will depend heavily on the type of bifurcation the neural network undergoes. If the bifurcation is supercritical, each input vector is mapped to a stable limit cycle with a location corresponding to one of the memorized pattern vectors. This is an example of a GAS-type neural network. On the other hand, if the bifurcation is subcritical, the trajectory corresponding to the initial input vector leaves the vicinity of the equilibrium point along a direction associated with one of the memorized patterns, which is an example of a NH-type neural network.
	
	Our earlier analysis of the coupled supercritical oscillators with normal connectivity matrix directly applies to this setting, since according to the Perron-Frobenuis theorem, a change in input corresponds to moving the leading eigenvalue as we assumed above in our calculations. Thus, the network parameter $||a||$, can switch the neural network type between GAS (supercritical) and NH (subcritical). 
	
	\section*{Conclusion}
	
	Utilizing center manifold reduction theory and Lyapunov coefficients, we have shown that a network of randomly coupled supercritical Andronov-Hopf oscillators can undergo either a supercritical or, surprisingly, a subcritical Andronov-Hopf bifurcation. We have also discussed how this switching between supercritical and subcritical regimes could potentially lead to a mechanism for neural network type switching in cortex. It is feasible that feedback connections from higher cortical areas could alter the activity of inhibitory interneurons in the network \cite{gilbert2007, gilbert2013}, leading to global changes in network parameters such as $||a||$. More experiments are needed to determine if the switching of dynamical regimes as presented in this paper might underlie any of the adaptive processing behaviors observed in intrinsic cortical circuitry.

\end{document}